OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 20 meters and a height of 30 meters? Answer 1,800 m3 4,000 m3 600 m3 12,000 m3

OpenStudy (saifoo.khan):

Did you tried using the formula?

OpenStudy (anonymous):

i dont understand this

OpenStudy (saifoo.khan):

No prob. Use the formula, \[\frac13 * \text{base area} \times height\]

OpenStudy (saifoo.khan):

\[\frac13 \times (20*20) \times (30)\]Now? :0

OpenStudy (anonymous):

so its B